OpenStudy (anonymous):

Solve by factoring: 3x2 + 2x - 1 = 0 A. {-1, 2/3} B. {1, -2/3} C. {-1, 1/3} D. {1, -1/3}

OpenStudy (precal):

c

OpenStudy (hoblos):

3x2 + 2x - 1 = 0 (3x-1)(x+1) =0 3x-1=0 or x+1=0 x=1/3 x=-1 {-1, 1/3}